OpenStudy (anonymous):

The sum of three numbers is 69. If the second number is equal to the first diminished by 8, and the third number is 5 times the first. What are the numbers? If x represents the first number, then which of the following equations could be used to solve the problem? x = 6x - 8 69 = 6x - 8 69 = 7x - 8 i got b

OpenStudy (anonymous):

x + (x - 8) + 5x = 69 2x + 5x - 8 = 69 7x - 8 = 69

OpenStudy (anonymous):

No "C" would be the correct answer.. :D

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

oops meant c